pooja195 (pooja195):

Evaluate 3x^3 when x=2 (this is a test(not a test test but a test :P )

OpenStudy (alexandervonhumboldt2):

are you kidding pooja? \[3*2^3\] 2^3=2*2*2=?

pooja195 (pooja195):

Is there something wrong with me asking a question?

OpenStudy (just_one_last_goodbye):

ITS A TEST :O

OpenStudy (just_one_last_goodbye):

smh tests

pooja195 (pooja195):

No lol they were supposed to but in a fature

pooja195 (pooja195):

*feature

OpenStudy (astrophysics):

\[x^3 \implies x \times x \times x\] the exponent tells you how many times you multiply by itself. \[3(2)^3 \implies 3[(2)(2)(2)] = ?\]

pooja195 (pooja195):

81

OpenStudy (justin528p):

24 bro

pooja195 (pooja195):

Thanks guys/gals :P

OpenStudy (shiburinga):

And I get suspended for asking questions on a test... ._.

pooja195 (pooja195):

3(3^3) 3^1*3^3 (3)*(3*3*3) 3*3*3*3 3^4 81

pooja195 (pooja195):

Not 24

OpenStudy (astrophysics):

I thought x = 2?

OpenStudy (astrophysics):

But in any case, you got the idea.

OpenStudy (justin528p):

2x2x2is 8 8x3 is 24

pooja195 (pooja195):

Whoops .-. lol thanks :P And its not a test >_>

OpenStudy (just_one_last_goodbye):

ITS A TEST :O

OpenStudy (shiburinga):

(It's a test but not a test) What the heck am I supposed to think? e.e

OpenStudy (just_one_last_goodbye):

actually i just want to bother Pooja xD shes my friend... so please dont involve me in your judgments

pooja195 (pooja195):

Thanks for all the help guys :)

OpenStudy (sneakysquash):

I Love Makonnen- Swerve

OpenStudy (alexandervonhumboldt2):

pooja joked i guess

pooja195 (pooja195):

No if you read what I wrote OpenStudy was supposed to add a feature but i guess they didnt. I wanted to see if they did.

OpenStudy (shiburinga):

Lol Poojy I was looking out for you like you look out fo rme :o

OpenStudy (alexandervonhumboldt2):

well opesntudy was guarding new featurws, so i guess pooja mde a fake question to test sometihng lol

OpenStudy (astrophysics):

A tip for next time, when you're testing something, don't say it's a test, as you will get spam as shown above which can be quite irritating, Ask your question and observe.

pooja195 (pooja195):

Btw when someone asks a question its best to just help them and not say things like "Are you kidding me" that can hurt someones feelings. (not mine) but others. @Astrophysics will do :P i didnt want people saying "are you stupid" <---i get that ALOT -.- for the questions i need legit help with.

OpenStudy (astrophysics):

Don't worry about what others say, if you have a question just ask, I don't judge and if others do they shouldn't be teaching.

pooja195 (pooja195):

:) i'll remember that